OpenStudy (anonymous):

The length of a triangle is 7 cm more than five times the width. The perimeter is 254cm . find the dimensions of the triangle. I need answer pleasee

OpenStudy (disco619):

 Perimeter = 254cm AB + BC + CA = 254cm

OpenStudy (disco619):

x + 7 + 5x + CA = 254

OpenStudy (disco619):

Suppose CA is "y" \[x + 7 + 5x + y = 254\] Now I think you should be able to solve this ahead...

OpenStudy (anonymous):

@Disco619 you have two variables you can not solve this ahead

OpenStudy (anonymous):

either the triangle has to be an isoceles

OpenStudy (anonymous):

@trix11 are u sure the question mentioned the length of a triangle,, because it should be of another shape...please check

OpenStudy (anonymous):

@mohammad.arqum Sorry . Its Rectangle.

OpenStudy (anonymous):

see now that makes the question loads easier

OpenStudy (disco619):

^that's what the equation comes considering if only "One" side of the triangle is "x". If it's an "Isosceles" triangle then simply do this:  Because an Isosceles triangle has two sides equal. \[x+x+7+5x=254\] \[2x+7+5x = 254\] \[7x + 7 = 254\] \[7x = 254 - 7\] \[7x = 247\] \[x = \frac{ 247 }{ 7 }\] \[x = 35.28571428571429\] Now if you have to round it off, do that and then Substitute it's value with the "x" in: \[7 + 5x\] to find the Length. And the value of "x" (width of triangle) itself would be "35.28571428571429"

OpenStudy (anonymous):

@Disco619 would u try to do it

OpenStudy (disco619):

LOL A RECTANGLE!

OpenStudy (anonymous):

LOL

OpenStudy (disco619):

Im done! @mohammad.arqum , you do it. [facepalm]

OpenStudy (anonymous):

okay length=x width=5x+7 2x+10x+14=254 12x=240 x=20 length=20 width=107

OpenStudy (anonymous):

Oh :D thank you so much the two of you . @mohammad.arqum @Disco619 big help :)

OpenStudy (anonymous):

No biggy @trix11